OpenStudy (anonymous):

Hellllllppppppp! Thank You :) What are the solutions of 3x2 + 15 = -6x? -1 ± i -1 ± 2i -2 ± i -2 ± 2i

OpenStudy (anonymous):

My Answer Is A

OpenStudy (anonymous):

@phi

OpenStudy (phi):

First put your equation into standard form 3x2 + 15 = -6x add 6x to both sides to get \[ 3x^2 + 6x + 15 = 0 \] notice that each term can be divided by 3. Do that to simplify. (in other words, divide both sides of the equation by 3) \[ x^2 +2x + 5 =0 \] now use quadratic formula see http://www.khanacademy.org/math/algebra/quadratics/quadratic_formula/v/using-the-quadratic-formula or "complete the square"

OpenStudy (phi):

Though a bit painful, you could check each of the choices and see which one works. You chose -1 ± i Let's check -1 + i Let x= -1 + i. x^2 = (-1+i)(-1+i)= 1 -2i + i^2 = 1 -2i-1 = -2i put those values in the simplified version of the equation \[ x^2+2x+5=0 \\ -2i +2(-1+i)+5 \\ -2i -2+2i +5 = -2+5 = 3\] 3 is not 0, so A is not the answer.